OpenStudy (yiwen):

A 3.0-uC positive point charge is located at the origin and a 2.0uC positive point charge is located at x=0.00m,y=1.0m. Find the coordinates of the point where the net electric field strength due to these charges is zero. i want to ask the solution of this question ? ans is :x=0.00m,y=0.55m

OpenStudy (irishboy123):

You really should draw it (though I know the mob app makes that impossible if you're on that :(( ) Use Coulomb's Law: the electric field due to a point charge is \(\vec E = \dfrac{kQ}{r^2} ~ \hat r\) With **positive** charges at \((0,0)\) and \((0,1)\), the fields will cancel each other out somewhere along the y - axis between those points, we know that from symmetry and the Coulomb's Law/ common sense. If that point is at \((0, y)\), we can say their magnitudes are the same at \((0, y)\), ie \( \dfrac{k(3)}{y^2} = \dfrac{k (2)}{(1-y)^2} \) And then you solve that. You will get 2 solutions, use the sensible one :)

OpenStudy (irishboy123):

To be uber-pernickety: \[\dfrac{k(3 \cdot 10^{-6})}{y^2} = \dfrac{k (2\cdot 10^{-6})}{(1-y)^2}\]